# revtex4-2Repair the float

### v3

@@ Abstract
Pair density Waves (PDW) are a inhomogeneous superconducting states whose Cooper pairs posses a finite momentum resulting in a oscillatory gap in space, even in the absence of an external magnetic field. There is growing evidence for the existence of PDW superconducting order in many strongly correlated materials particularly in the cuprate superconductors and in several other different types of systems. A feature of the PDW state is that inherently it has a CDW as a composite order associated with it. Here we study the structure of the electronic topological defects of the PDW, paying special attention to the half-vortex and its electronic structure that can be detected in STM experiments. We discuss tell-tale signatures of the defects in violations of inversion symmetry, in the excitation spectrum and their spectral functions in the presence of topological defects. We discuss the “Fermi surface” topology of Bogoliubov quasi-particle of the PDW phases and we briefly discuss the role of quasi-particle-interference.
@@ 1 Introduction
One of the central problems in Condensed Matter Physics is the understanding the phases of strongly correlated systems such as high Tc superconductors. Experimental and theoretical research during the past decade has clearly shown that a characteristic feature of these systems is that their complex phase diagrams have phases with different types of charge, spin and superconducting orders which are intertwined rather than competing with each other [1]. The prototype quantum materials with intertwined orders are the cuprate superconductors, which possess a very rich phase diagram hosting antiferromagnetic order, high temperature d-wave superconductivity, CDW order, nematic order and, at least in the lanthanum family of cuprates, incommensurate SDW order. Experiments done during the past decade have provided increasing evidence that, in addition to the d-wave superconducting order, in these systems a novel type of superconducting order, known as a pair-density-wave may be at play [2]. In addition, and in contrast to conventional superconductors where the superconducting (SC) state is born from a normal Fermi liquid, the “normal” (high temperature) state of all the cuprates is a strange metal, a metal without well-defined electronic quasi-particle.
Of particular interest is the cuprate material La2-xBaxCuO2 (LBCO). This cuprate superconductor, the original high Tc material, has the interesting phase diagram provided in figure 1 [3]. Instead of a single SC dome, as most other cuprates have, La2-xBaxCuO4 has a pronounced anomaly at $x=1/8$ hole doping where the transition temperature to the d-wave SC state is suppressed dramatically from 35K to about 4K where the Meissner state is observed. In this regime a remarkable set of phase transitions are observed [4]: static charge-density-wave (CDW) and spin-density-wave (SDW) orders onset at  52K and  42K, respectively. Below the onset of the SDW order the c-axis resistivity, $\rho_{c}$, increases with decreasing temperature while the ab plane resistivity, $\rho_{ab}$, decreases rapidly. Superconducting phase fluctuations in the ab planes onset at about 35K and at approximately 16 K a two-dimensional Berezhinskii-Kosterlitz-Thouless (BKT) transition to a two dimensional SC state is observed. On the other hand, the resistive transition where the c-axis resistivity vanishes only happens at 10K and the full Meissner d-wave SC state is only reached at $T_{c}\sim 4$K. This “dynamical layer decoupling” behavior is also observed in La2-xBaxCuO4 in the presence of a c-axis magnetic field away from $x=1/8$ [5] and in underdoped La2-xSrxCuO4 in magnetic fields $B\sim~{}8$T [6] where a magnetic-field induced SDW was observed long ago [7]. Similar behaviors have been found in LSCO doped with Zn [8] and with iron [9].
@@
Evidence for PDW order also exists in cuprate superconductors that do not have the LTT crystal structure such as Bi2Sr2CaCu2O8+δ. In this case much of the evidence was found in STM experiments [13] and Josephson tunneling spectroscopy [14]. STM experiments in the vortex halo of Bi2Sr2CaCu2O8+δ have revealed tell-tale evidence for PDW order in that regime [15, 16]. Recent experiments have provided evidence for PDW order to be present also in the superconducting states of several novel strongly correlated materials including the heavy fermion material UTe2 [17, 18], the iron superconductor Fe(Se,Te) [19], the pnictide EuRbFe4As4 [20], and in kagome superconductors such as CsV3Sb5 [21, 22]. In this study we will focus on the pristine PDW with no uniform component present. In a future publication we will address how the uniform component affects various plots seen in this paper.
The PDW is a superconducting state in which Cooper pairs with finite center of mass momentum $\bm{Q}$ condense. In such a state the local pairing amplitude $\Delta(\bm{r})$ is periodic function of position whose period is $2\pi/|\bm{Q}|$. The order parameter of the PDW has the same symmetry as the Larkin-Ovchinnikov (LO) state [23]. However, the PDW differs from the LO state in several important ways: a) the LO state is created by a magnetic field through the Zeeman coupling to the spins and it is spin-polarized, and b) as a result, it has a broken time-reversal invariance. In contrast, the PDW arises in the absence of a magnetic field. Another superconducting state with a finite-momentum Cooper pair is the Fulde-Ferrell (FF) state [24] which, like the LO state, was also envisioned as arising in the presence of a Zeeman coupling to an external magnetic field. Much like the LO state, the FF state breaks time reversal but, in addition, it breaks inversion symmetry. The amplitude of the FF state is constant in space. Finally the ordering wave vectors of the LO and the FF states are tuned by the strength of the magnetic field and, consequently, the periodicity of both states are much larger than the lattice constant.
@@
Figure 1: (Online Color) Experimental data of LBCO indicating various phases for Temperature, $T$, vs hole-doping, $x$. Various orders exist simultaneously under the superconducting dome which may indicate the onset of the PDW phase. Data from Ref.[3].
@@
The phenomenology of the PDW state [10, 25, 26, 27, 2] (and the phenomenology of all high temperature superconductors) strongly suggests that, with some possible exceptions, a PDW is necessarily a strong coupling state which cannot be explained in terms of the conventional BCS theory of superconductivity [28, 29]. Nevertheless, BCS-type approaches have been developed to explain the PDW [30, 31, 32, 33] which require that the interactions be large compared to the kinetic energy of the holes. In this regime BCS theory is not reliable. Numerical simulations of $t-J$ and extended Hubbard models have been shown evidence that PDW ground is at least a strong competitor to be ground state  [34, 35, 36, 37, 38]. Quasi-one-dimensional models have also shown that their phase diagrams contain PDW phases [39]. Under special circumstances weak coupling models do predict the existence of PDW phases but typically they require some other strong coupling physics to take place first, such as a Pomeranchuk instability in the triplet channel [40]. We should note that a recent study [41] predicts the occurrence of a PDW state in transition metal dichalcogenide materials. To the best of our knowledge the only microscopic model which is unambiguously known to have a (large) PDW phase is the one-dimensional Kondo-Heisenberg chain [42, 43].
@@
The purpose of this paper is to investigate physics of the PDW state by studying the electronic structure of the topological defects of this state: the half-vortex, the double dislocation and the Abrikosov vortex. The study that we undertake here is relevant to the understanding of the features of these defects revealed by STM experiments. Here we adopt a phenomenological description of the PDW and we will not concern ourselves with the possible physical mechanism(s) associated with this superconducting state. For concreteness we will consider a system with a square lattice (kagome and honeycomb lattices have also been considered). On a square lattice the PDW state may be unidirectional, which breaks spontaneously both translation symmetry and the point group symmetry of the square lattice, or bidirectional, which is invariant under the point group symmetry but breaks translation symmetry along two directions. Let $\Delta(\bm{r})$ be the local amplitude for a spin singlet superconductor (where $\bm{r}$ is a lattice site; in the case of a local d-wave state $\bm{r}$ is the superconducting amplitude on a bond between two nearest neighbor sites $\bm{r}$ and $\bm{r}^{\prime}$). We will consider the simpler case of an unidirectional spin-singlet PDW with a period close to $8a_{0}$ which is appropriate for the lanthanum cuprates [2]. The local pairing amplitude can be expanded in Fourier components
$\Delta(\bm{r})=\Delta_{0}(\bm{r})+\Delta_{\bm{Q}}(\bm{r})e^{i\bm{Q}\cdot\bm{r}}+\Delta_{-\bm{Q}}(\bm{r})e^{-i\bm{Q}\cdot\bm{r}}$ (1.1)
@@
In this paper we will consider only a unidirectional PDW state and hence assume that $\langle\Delta_{0}\rangle=0$. The central results of this paper is concerned with experimental signatures associated with the induced CDW “daughter” states of the PDW parent state [25, 45]. The induced CDW order with wave vector $\bm{K}=2\bm{Q}$ whose order parameter field is $\rho_{\bm{2}\bm{Q}}(\bm{r})\sim\Delta_{\bm{Q}}\Delta_{-\bm{Q}}^{*}$ (and similarly for $\rho_{-2\bm{Q}}(\bm{r})$) is described in more detail in section 2, including the topological defects of the PDW state we are interested in. The unidirectional PDW has two complex order parameters, $\Delta_{\pm\bm{Q}}(\bm{r})$, which means two amplitude fields and two phase fields. This means that the order parameter of the PDW transforms under a $U(1)\times U(1)$ global symmetry, where the first factor is the conventional global $U(1)$ gauge invariance of a superconductor and the second factor represents the invariance under continuous translations of the incommensurate PDW state. In Refs. [25, 45] it is shown that as a result of the topology of the target space of the order parameter(s) of the PDW, this state has three distinct topological defects: a superconducting Abrikosov vortex, a half-vortex, and a double dislocation.
@@
The electronic structure of the PDW has Bogoliubov quasi-particle with (Bogoliubov) Fermi surfaces which define pockets of quasi-particle states [25, 46]. The stability of Bogoliubov Fermi surfaces has been established in several studies [47, 48]. In the PDW state the half-vortex has an energy cost which diverges logarithmically with sample size. In contrast, in the presence of a non-vanishing uniform SC component, $\langle\Delta_{0}\rangle\neq 0$, the energy cost is instead linearly divergent and, hence, half-vortices and anti-half-vortices are confined into pairs. Experimental evidence for half-vortices pairs has been found in STM experiments in Bi2Sr2CaCu2O8+δ by Du and coworkers [14].
@@
In this work we describe the half-vortex as a state in which one of the two order parameters of the PDW, say $\Delta_{\bm{Q}}(\bm{r})$, has a vortex while the other order parameter, $\Delta_{-\bm{Q}}(\bm{r})$, does not. Since one of the SC order parameters vanishes at the defect location while the other does not, at the defect core we have an FF-type state which breaks inversion symmetry. In this sense, the half-vortex has an FF halo. This behavior is analogous to the halo associated with the Abrikosov vortex in a system in which uniform SC and PDW orders compete [15, 16]. Our construction of the half-vortex follows the same strategy used by Wang and coworkers [16] for the Abrikosov vortex halo of an uniform superconductor with a PDW as a subleading order. In contrast, in the case of a Abrikosov vortex both PDW order parameters wind with the same topological charge whereas in the double dislocation they wind with opposite topological charges.
@@
We then investigate the electronic structure of the PDW in the presence of topological defects by embedding the resulting configuration into the Bogoliubov-de Gennes Hamiltonian. Here we use a non-interacting band structure suitable for a superconductor in a copper oxide plane. Using this effective Hamiltonian with the defect background we calculated the local density-of-states and a function of bias for a model of an STM measurement at a point contact with a normal metal. We also compare the results on the half-vortex with the same calculation done for a full Abrikosov vortex and for a double dislocation of the CDW.
@@
The paper is organzied as follows. In section 2 we introduce the Landau-Ginzburg equations to be used in our analysis. In this section the notion of intertwined order is reviewed and we discuss the composite order parameters of interest in subsection 2.2. An overview of the relationship between experiment and induced order is briefly discussed there as well. In subsection 2.3 we review the topological defects of the PDW superconducting state. Section 3 is devoted to the construction of a static half-vortex. Here we discuss how the profiles of the components of the PDW order parameters are modeled in the case of a half-vortex. In section 4 the effective mean field Hamiltonian for the PDW with static topological defects used in our simulation is introduced as well as the Green functions and the related spectral function. Here we present results for the local-density-of-states (LDOS), used to acquire plots for the numerical solutions. In section 4 the numerical parameters and the explicit form of the order parameters used in the simulation are discussed. In section 4.2 the plots of the charge-density wave order of the PDW in the presence of defects are shown, and various other aspects of the vortex cores are explored/discussed. In section 4.5 we discuss the spectral functions of the PDW order and display the dispersion relation for a PDW with a half-vortex. Finally, in section 5 we discuss some implications of our results and summarize the most salient results. Several appendices are devoted to technical details. Details of the Bogoliubov transformation are given in appendix A and the setup for the numerical diagonalization of the Bogoliubov-de Gennes Hamiltonian is sketched in appendix B. In appendix C we give details of the retarded Green function at zero temperature. In appendix E we compare the tunneling DOS for a PDW, an FF state and for the half-vortex of the PDW. In appendix F we present data on spectral functions for PDW states with s and d wave form factors.

### v4

@@ Abstract
Pair density waves (PDWs) are a inhomogeneous superconducting states whose Cooper pairs possess a finite momentum resulting in a oscillatory gap in space, even in the absence of an external magnetic field. There is growing evidence for the existence of PDW superconducting order in many strongly correlated materials, particularly in the cuprate superconductors and in several other different types of systems. A feature of the PDW state is that inherently it has a CDW as a composite order associated with it. Here we study the structure of the electronic topological defects of the PDW, paying special attention to the half-vortex and its electronic structure that can be detected in STM experiments. We discuss tell-tale signatures of the defects in violations of inversion symmetry, in the excitation spectrum and their spectral functions in the presence of topological defects. We discuss the “Fermi surface” topology of Bogoliubov quasiparticle of the PDW phases, and we briefly discuss the role of quasiparticle interference.
@@ 1 Introduction
One of the central problems in condensed matter physics is understanding the phases of strongly correlated systems such as high-Tc superconductors. Experimental and theoretical research during the past decade has clearly shown that a characteristic feature of these systems is that their complex phase diagrams have phases with different types of charge, spin and, superconducting orders which are intertwined rather than competing with each other [1]. The prototype quantum materials with intertwined orders are the cuprate superconductors, which possess a very rich phase diagram hosting antiferromagnetic order, high-temperature $d$-wave superconductivity, CDW order, nematic order, and, at least in the lanthanum family of cuprates, incommensurate SDW order. Experiments done during the past decade have provided increasing evidence that, in addition to the $d$-wave superconducting order, in these systems a type of superconducting order, known as a pair density wave, may be at play [2]. In addition, and in contrast to conventional superconductors where the superconducting (SC) state is born from a normal Fermi liquid, the “normal” (high-temperature) state of all the cuprates is a strange metal, a metal without well-defined electronic quasiparticle.
Of particular interest is the cuprate material La2-xBaxCuO2 (LBCO). This cuprate superconductor, the original high Tc material, has the interesting phase diagram provided in Fig. 1 [3]. Instead of a single SC dome, as most other cuprates have, La2-xBaxCuO4 has a pronounced anomaly at $x=1/8$ hole doping where the transition temperature to the $d$-wave SC state is suppressed dramatically from 35K to about 4K where the Meissner state is observed. In this regime a remarkable set of phase transitions are observed [4]: static charge density-wave (CDW) and spin-density-wave (SDW) orders onset at  52K and  42K, respectively. Below the onset of the SDW order the $c$-axis resistivity, $\rho_{c}$, increases with decreasing temperature while the $ab$-plane resistivity, $\rho_{ab}$, decreases rapidly. Superconducting phase fluctuations in the $ab$-planes onset at about 35K, and at approximately 16 K a two-dimensional Berezhinskii-Kosterlitz-Thouless (BKT) transition to a two-dimensional SC state is observed. On the other hand, the resistive transition where the $c$-axis resistivity vanishes happens only at 10K, and the full Meissner $d$-wave SC state is reached only at $T_{c}\sim 4$K. This “dynamical layer decoupling” behavior is also observed in La2-xBaxCuO4 in the presence of a $c$-axis magnetic field away from $x=1/8$ [5] and in underdoped La2-xSrxCuO4 in magnetic fields $B\sim~{}8$T [6] where a magnetic-field induced SDW was observed long ago [7]. Similar behaviors have been found in LSCO doped with Zn [8] and with iron [9].
@@
Evidence for PDW order also exists in cuprate superconductors that do not have the LTT crystal structure such as Bi2Sr2CaCu2O8+δ. In this case much of the evidence was found in STM experiments [13] and Josephson tunneling spectroscopy [14]. STM experiments in the vortex halo of Bi2Sr2CaCu2O8+δ have revealed tell-tale evidence for PDW order in that regime [15, 16]. Recent experiments have provided evidence for PDW order to be present also in the superconducting states of several strongly correlated materials including the heavy fermion material UTe2 [17, 18], the iron superconductor Fe(Se,Te) [19], the pnictide EuRbFe4As4 [20], and kagome superconductors such as CsV3Sb5 [21, 22]. In this study we will focus on the pristine PDW with no uniform component present. In a future publication we will address how the uniform component affects various plots seen in this paper.
The PDW is a superconducting state in which Cooper pairs with finite center of mass momentum $\bm{Q}$ condense. In such a state the local pairing amplitude $\Delta(\bm{r})$ is periodic function of position whose period is $2\pi/|\bm{Q}|$. The order parameter of the PDW has the same symmetry as the Larkin-Ovchinnikov (LO) state [23]. However, the PDW differs from the LO state in several important ways: (1) the LO state is created by a magnetic field through the Zeeman coupling to the spins and it is spin-polarized, and (2) as a result, it has a broken time-reversal invariance. In contrast, the PDW arises in the absence of a magnetic field. Another superconducting state with a finite-momentum Cooper pair is the Fulde-Ferrell (FF) state [24], which, like the LO state, was also envisioned as arising in the presence of a Zeeman coupling to an external magnetic field. Much like the LO state, the FF state breaks time reversal, but, in addition, it breaks inversion symmetry. The amplitude of the FF state is constant in space. Finally the ordering wave vectors of the LO and the FF states are tuned by the strength of the magnetic field, and, consequently, the periodicity of both states is much larger than the lattice constant.
@@
Figure 1: (Online Color) Experimental data of LBCO indicating various phases for temperature, $T$, vs hole-doping, $x$. Various orders exist simultaneously under the superconducting dome, which may indicate the onset of the PDW phase. Data from Ref. [3].
@@
The phenomenology of the PDW state [10, 25, 26, 27, 2] (and the phenomenology of all high-temperature superconductors) strongly suggests that, with some possible exceptions, a PDW is necessarily a strong coupling state, which cannot be explained in terms of the conventional BCS theory of superconductivity [28, 29]. Nevertheless, BCS-type approaches have been developed to explain the PDW [30, 31, 32, 33] which require that the interactions be large compared to the kinetic energy of the holes. In this regime BCS theory is not reliable. Numerical simulations of $t-J$ and extended Hubbard models have shown evidence that PDW ground is at least a strong competitor to be ground state  [34, 35, 36, 37, 38]. Quasi-one-dimensional models have also shown that their phase diagrams contain PDW phases [39]. Under special circumstances weak coupling models do predict the existence of PDW phases, but typically they require some other strong coupling physics to take place first, such as a Pomeranchuk instability in the triplet channel [40]. We should note that a recent study [41] predicts the occurrence of a PDW state in transition metal dichalcogenide materials. To the best of our knowledge the only microscopic model which is unambiguously known to have a (large) PDW phase is the one-dimensional Kondo-Heisenberg chain [42, 43].
@@
The purpose of this paper is to investigate physics of the PDW state by studying the electronic structure of the topological defects of this state: the half-vortex, the double dislocation and the Abrikosov vortex. The study that we undertake here is relevant to the understanding of the features of these defects revealed by STM experiments. Here we adopt a phenomenological description of the PDW and we will not concern ourselves with the possible physical mechanism(s) associated with this superconducting state. For concreteness we will consider a system with a square lattice (kagome and honeycomb lattices have also been considered). On a square lattice the PDW state may be unidirectional, which breaks spontaneously both translation symmetry and the point group symmetry of the square lattice, or bidirectional, which is invariant under the point group symmetry but breaks translation symmetry along two directions. Let $\Delta(\bm{r})$ be the local amplitude for a spin singlet superconductor (where $\bm{r}$ is a lattice site; in the case of a local $d$-wave state $\bm{r}$ is the superconducting amplitude on a bond between two nearest-neighbor sites $\bm{r}$ and $\bm{r}^{\prime}$). We will consider the simpler case of an unidirectional spin-singlet PDW with a period close to $8a_{0}$, which is appropriate for the lanthanum cuprates [2]. The local pairing amplitude can be expanded in Fourier components
$\Delta(\bm{r})=\Delta_{0}(\bm{r})+\Delta_{\bm{Q}}(\bm{r})e^{i\bm{Q}\cdot\bm{r}}+\Delta_{-\bm{Q}}(\bm{r})e^{-i\bm{Q}\cdot\bm{r}},$ (1.1)
@@
In this paper we will consider only a unidirectional PDW state and hence assume that $\langle\Delta_{0}\rangle=0$. The central results of this paper are concerned with experimental signatures associated with the induced CDW “daughter” states of the PDW parent state [25, 45]. The induced CDW order with wave vector $\bm{K}=2\bm{Q}$ whose order parameter field is $\rho_{\bm{2}\bm{Q}}(\bm{r})\sim\Delta_{\bm{Q}}\Delta_{-\bm{Q}}^{*}$ [and similarly for $\rho_{-2\bm{Q}}(\bm{r})$] is described in more detail in Sec. 2, including the topological defects of the PDW state we are interested in. The unidirectional PDW has two complex order parameters, $\Delta_{\pm\bm{Q}}(\bm{r})$, which means two amplitude fields and two phase fields. This means that the order parameter of the PDW transforms under a $U(1)\times U(1)$ global symmetry, where the first factor is the conventional global $U(1)$ gauge invariance of a superconductor and the second factor represents the invariance under continuous translations of the incommensurate PDW state. In Refs. [25, 45] it is shown that as a result of the topology of the target space of the order parameter(s) of the PDW, this state has three distinct topological defects: a superconducting Abrikosov vortex, a half-vortex, and a double dislocation.
@@
The electronic structure of the PDW has Bogoliubov quasiparticle with (Bogoliubov) Fermi surfaces which define pockets of quasiparticle states [25, 46]. The stability of Bogoliubov Fermi surfaces has been established in several studies [47, 48]. In the PDW state the half-vortex has an energy cost which diverges logarithmically with sample size. In contrast, in the presence of a non-vanishing uniform SC component, $\langle\Delta_{0}\rangle\neq 0$, the energy cost is instead linearly divergent and, hence, half-vortices and anti-half-vortices are confined into pairs. Experimental evidence for half-vortices pairs has been found in STM experiments in Bi2Sr2CaCu2O8+δ by Du and coworkers [14].
@@
In this work we describe the half-vortex as a state in which one of the two order parameters of the PDW, say $\Delta_{\bm{Q}}(\bm{r})$, has a vortex while the other order parameter, $\Delta_{-\bm{Q}}(\bm{r})$, does not. Since one of the SC order parameters vanishes at the defect location while the other does not, at the defect core we have an FF type state which breaks inversion symmetry. In this sense, the half-vortex has an FF halo. This behavior is analogous to the halo associated with the Abrikosov vortex in a system in which uniform SC and PDW orders compete [15, 16]. Our construction of the half-vortex follows the same strategy used by Wang and coworkers [16] for the Abrikosov vortex halo of an uniform superconductor with a PDW as a subleading order. In contrast, in the case of a Abrikosov vortex both PDW order parameters wind with the same topological charge, whereas in the double dislocation they wind with opposite topological charges.
@@
We then investigate the electronic structure of the PDW in the presence of topological defects by embedding the resulting configuration into the Bogoliubov-de Gennes Hamiltonian. Here we use a non-interacting band structure suitable for a superconductor in a copper oxide plane. Using this effective Hamiltonian with the defect background we calculated the local density of states and a function of bias for a model of an STM measurement at a point contact with a normal metal. We also compare the results on the half-vortex with the same calculation done for a full Abrikosov vortex and for a double dislocation of the CDW.
@@
The paper is organized as follows. In Sec. 2 we introduce the Landau-Ginzburg equations to be used in our analysis. In this section the notion of intertwined order is reviewed and we discuss the composite order parameters of interest in Sec. 2.2. An overview of the relationship between experiment and induced order is briefly discussed there as well. In Sec. 2.3 we review the topological defects of the PDW superconducting state. Section 3 is devoted to the construction of a static half-vortex. Here we discuss how the profiles of the components of the PDW order parameters are modeled in the case of a half-vortex. In Sec. 4 the effective mean field Hamiltonian for the PDW with static topological defects used in our simulation is introduced as well as the Green functions and the related spectral function. Here we present results for the local density of states (LDOS), used to acquire plots for the numerical solutions. In Sec. 4 the numerical parameters and the explicit form of the order parameters used in the simulation are discussed. In Sec. 4.2 the plots of the charge density wave order of the PDW in the presence of defects are shown, and various other aspects of the vortex cores are explored and discussed. In Sec. 4.5 we discuss the spectral functions of the PDW order and display the dispersion relation for a PDW with a half-vortex. Finally, in Sec. 5 we discuss some implications of our results and summarize the most salient results. Several appendixes are devoted to technical details. Details of the Bogoliubov transformation are given in Appendix A and the setup for the numerical diagonalization of the Bogoliubov-de Gennes Hamiltonian is sketched in Appendix B. In Appendix C, we give details of the retarded Green function at zero temperature. In Appendix E we compare the tunneling DOS for a PDW, an FF state and for the half-vortex of the PDW. In Appendix F we present data on spectral functions for PDW states with $s$- and $d$- wave form factors.